AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME, Page 12 of 33
New
|
Find
|
Select
|
Import
|
Config
|
Last day
|
Help
Full
| Summary |
Threaded
-- All entries --
Last day
Last 3 Days
Last week
Last month
Last 3 Months
Last 6 Months
Last Year
655 Entries
Goto page
Previous
1
,
2
,
3
...
11
, 12,
13
...
31
,
32
,
33
Next
ID
Date
Author
Subject
Text
440
Sat Sep 2 10:04:55 2023
CB
Turbo GV closed
Found CARME in XHV at 1.1E-11 mbar. Chamber
temperature around 30C cooling slowly.
Closed GV to top turbo. Pressure
439
Fri Sep 1 08:11:40 2023
CB
NEG activation - 2
9:10 Set heaters temperature 300 ->
250 C in preparation for cooling down and
to help keep DSSD temperature down. Started
438
Thu Aug 31 08:36:34 2023
CB
NEG activation
Found CARME at 50C, a bit cooler than expected,
probably due to invasive tent changes yesterday
to mount missing wires. Increased heaters
437
Wed Aug 30 10:28:36 2023
CB
Preparing for pumps on
We need to connect the non-bakeable (max
80C) wire of the ion pump belonging to the
vacuum division, and we also need to connect
436
Wed Aug 30 09:33:13 2023
CB, OG
NEG cables repaired
Repaired Neg wires for D2000 with Oemer's
help.
Cable #2 is fine.
435
Mon Aug 14 09:22:56 2023
JM
Bakeout
14/08/23
Temperature in the baking volume
is ~ 42 C and the pressure is 4.6e-8 mbar.
434
Sat Aug 12 15:06:51 2023
JM
Initial bakeout tests
12/08/23
Baking testing. Aim to find relationship
between heater temperature setpoint and the
433
Sat Aug 12 11:57:51 2023
JM
Updated thermocouple positions and heater channel layouts
Heater channel layouts for the interaction
chamber and carme in attached documents.
Updated thermocouple positions;
6x
431
Thu Aug 10 07:51:43 2023
JM
Baking tent and heater testing
09/08/23
Attached is the current pumpdown
plot taken from grafana.
8x
430
Thu Jul 27 09:43:56 2023
RSS
Thermocouples connected -2
10:30 IE514 1.80e-08mbar
Most
of the CARME is covered by the baking tent.
The tent needs to be closed and a bit covered
429
Wed Jul 26 08:28:09 2023
RSS
Building tent
09:26 IE514 1.96e-08mbar
428
Tue Jul 25 09:12:15 2023
RSS
Building tent
10:10 IE514 2.0e-08mbar
Upstream
part of the CARME is covered.
427
Mon Jul 24 10:40:14 2023
RSS, TD
Pump down continues
11.40 Pirani 1.00e-3 mbar IE514 2.23e-08mbar
15.15 7d/24h vacuum plots
426
Sun Jul 23 10:13:48 2023
RSS, TD
Building tent
11.15 Pirani 1.0e-03mbar IE514 2.72e-08mbar
16.35 Pirani 1.0e-03mbar IE514 2.61e-08mbar
425
Sat Jul 22 09:19:59 2023
RSS, TD
Thermocouples connected
10.20 Pirani 1e-03mbar IE514 3.26e-08mbar
10x
424
Fri Jul 21 08:21:54 2023
TD, RSS
Baking tent manual
09.20 Pirani 1.00e-3 mbar IE514 3.90e-8 mbar
Today - install TCs, connect to Pico
423
Thu Jul 20 08:29:26 2023
TD
Installing tent. Two NEG power cable broken.
09.28 Pirani 1.00e-3 mbar IE514 5.45e-8 mbar
12.10 Bakeout tent frame installed
11x
422
Wed Jul 19 08:17:26 2023
TD
Leak testing
09.16 Pirani 1.00e-3 mbar IE514 8.77e-8 mbar
Leak tester c. 5e-12 mbar.l/s (leak
tester valved off)
421
Tue Jul 18 08:16:28 2023
TD
Pumping down. Issues found with one detector.
09.15 Following overnight pumping of CARME
chamber by Leybold turbo
Pirani 1.00e-3 mbar IE514 2.83e-7 mbar
420
Mon Jul 17 08:45:11 2023
TD
Closing CARME again
Grosmantage are scheduled to arrive 10.00
to inspect COF600 seals and their fit to
Goto page
Previous
1
,
2
,
3
...
11
, 12,
13
...
31
,
32
,
33
Next
ELOG V3.1.4-unknown